A plot of land close to the beach could earn its owners $300/day as a parking lot or $500/day as a campground. Calculate the weekly economic rent of using the land as a campground.
Answer #1

Ans) $1,400

weekly economic rent of using the land as a campground.

weekly economic rent = 7*500 - 7*300 = 3,500 - 2,100 = 1,400